A 2.000 g piece of copper was used to precipitate silver ions aa silver solution. When the copper metal reacts, Cu2+ ions are formed which are soluble in water. If 0.750 g of Ag(s) were produced in this process what should the final mass of the copper wire be?
1
Expert's answer
2020-09-14T08:23:53-0400

QUESTION # 132910

A 2.000 g piece of copper was used to precipitate silver ions aa silver solution. When the copper metal reacts, Cu2+ ions are formed which are soluble in water. If 0.750 g of Ag(s) were produced in this process what should the final mass of the copper wire be?

ANSWER

Equation for the reaction

Cu(s) +2Ag+(aq)       ͢     Cu+2(aq) + 2Ag(s)

Moles of Ag formed =0.750/107.868

                                    = 0.006953 mol

Moles of Cu used = ½ (0.006953) mol

                               = 0.0034765 mol

Mass of Cu used = 0.0034765×63.546

                              = 0.2209g

Final mass of Cu wire = 2.000-0.2209

                                      = 1.779g